testing for nitrate analysis?
1
Expert's answer
2016-11-08T14:43:11-0500
A nitrate test is a chemical test used to determine the presence of nitrate ion in the object. A common nitrate test (brown ring test) can be performed by adding iron (II) sulfate to a solution of a nitrate, then slowly adding concentrated sulfuric acid such that the acid forms a layer below the aqueous solution.

As a result, a brown ring is formed at the junction of the two layers, indicating the presence of the nitrate ion. The presence of nitrite ions will interfere with this test.

Chemical process undergoing during this test are:

2HNO3+ 3H2SO4 + 6FeSO4 = 3Fe2(SO4)3 + 2NO + 4H2O

FeSO4 + 6H2O = [Fe(H2O)6]SO4

[Fe(H2O)6]SO4 + NO = [Fe(H2O)5(NO)]SO4+ H2O